OpenStudy (anonymous):

solve the equation 13x - 28 = 9x + 56

OpenStudy (lgbasallote):

subtract 9x from both sides...what do you get?

OpenStudy (anonymous):

19

OpenStudy (epikrebel):

The answer is not 19.

OpenStudy (anonymous):

28 - 9 is 19

OpenStudy (epikrebel):

You combine or subtract like terms. \[13x(-9x)-28(+28)=9x(-9x)+56(+28)\]

OpenStudy (epikrebel):

\[4x=84\] \[\frac{ 4x }{ 4 }=\frac{84}{4}\] x=21